OpenStudy (anonymous):

Find the general solution to the following sin(theta)cos2(theta)+sin(theta)=1+cos2(theta)

OpenStudy (raden):

sinxcos2x + sinx = 1+cos2x sinx(1+cos2x) = 1+cos2x sinx(1+cos2x) - (1+cos2x) = 0 (1+cos2x)(sinx-1) = 0 for zeros, take 1+cos2x=0 or sinx-1=0 * 1+cos2x=0 cos2x=-1 cos2x=cos(k*2pi +- pi) 2x=k*2pi +- pi x = k*pi +- pi/2 ** sinx-1=0 sinx = 1 sinx = sin pi/2 sinx =sin(k*2pi + pi/2) x =k*2pi + pi/2

OpenStudy (anonymous):

thank you so much

OpenStudy (raden):

welcome

OpenStudy (anonymous):

i have a question on the first problem why is it a negative instead of a positive if cos(pi)=-1?